how many three digit numbers can be formed from the digits 1,2,3,4,and 5 if each digit can be used only once
1
Expert's answer
2012-11-20T09:07:20-0500
Solution

n=5; k=3; A=n*(n-1)(n-2)...(n-(k-1))=5·4·3=60

Answer: 60 three digit numbers